# Non-Markovian Dynamics in Fiber Delay-Line Buffers

### v2

@@ II Results
We characterize the prepared state $|\psi_{\circ}\rangle=(|H_{s}H_{i}\rangle+|V_{s}V_{i}\rangle)/\sqrt{2}$ by using the standard method of QST James_pra_2001 ; White_pra_2001 ; Munro_pra_2001 . There are a total of 16 settings for the HWPs and QWPs in the $\rm{PA_{s}}$ and $\rm{PA_{i}}$. For each setting, we measure the coincidence counts (CC) and accidentals (AC) with an integration time of 2 seconds or 100M sampling gates when during the switches are turned on. Using the maximum likelihood estimation, we reconstruct the $4\times 4$ density matrix of the final state $\rho^{f}$ in the HV basis. We construct the density matrix $\rho^{f}$ based on the true coincidence counts i.e. the accidental coincidence is subtracted. In other words, the Raman photon and loss-induced accidental do not contribute to this study. We repeat the state tomography for different fiber lengths from 0.0 m up to 25 km. The 0.0 m fiber length corresponds to the short path in the buffer.
@@
We use the quantum discord to characterize the non-Markovian dynamics of the buffered photon. It is of both fundamental and practical interest to differentiate classical and quantum correlations within the total correlation for the buffered photon experiencing a non-Markovian environment. Two scenarios can contribute to the classical information; (1) environment projects or makes a measurement on the idler qubit state and creates the classical state, (2) environment perturbation can dechore the information of the idler qubit state such as degradation of the polarization properties. Scenario (1) is the classical correlation which will appear at the coincidence and accidental coincidence. We can avoid this classical correlation by subtracting the accidental. Scenario (2) can create a mixture of product states. We study the quantum discord under scenario (2) which can allow us to have a buffer time $>$ 0.9 ms. For this purpose, we can accurately treat the final state as the Werner state. The Werner state is the mixture of the maximally entangled state with the probability $P$ and the maximally mixed state as given by Werner_pra_1989 ; Munro_pra_2001 ,
@@
where the $I$ is the identity matrix. We calculate the average of the probability $P=\sum^{n=6}_{i}p^{\prime}_{i}/n$, where $p^{\prime}_{1(2)}=4\rho^{\prime}_{11(44)}-1$, $p^{\prime}_{3(4)}=2\rho^{\prime}_{14(41)}$, and $p^{\prime}_{5(6)}=1-4\rho^{\prime}_{22(33)}$. The density matrix elements $\rho^{\prime}_{11}$, $\rho^{\prime}_{44}$, $\rho^{\prime}_{14}$, $\rho^{\prime}_{41}$, $\rho^{\prime}_{22}$, and $\rho^{\prime}_{33}$ are obtained from the final state. The measured average probability $P$ is the non-Markovian probability function p(t). ( See Method for Amplitude damping on one photon from two-photon Werner state).
@@
We plot the average probability $P$ of the Werner state as a function of buffer time $t$ as shown (red solid circle) in Fig.3. We fit the experimental data with the summation of the probability functions $\mathcal{P}_{a}$ of Eq.(3) and $\mathcal{P}_{sy}$ of Eq.(4). With the substitution of $L=(c/n_{r})t$ and $\sqrt{\langle\tau^{2}\rangle}=D_{p}\,\sqrt{L}$, the $\Delta\phi_{1,2}$ in the harmonic functions of $\mathcal{P}_{a}$ and $\mathcal{P}_{sy}$ is then given by a $\Delta\phi_{1,2}=\Delta\omega\,\sqrt{\langle\tau_{1,2}^{2}\rangle}=\Delta\omega\,D_{p1,p2}\,\sqrt{\frac{c\cdot t}{n_{r}}}$, respectively, where $D_{p1,p2}$ is the PMD parameter in unit of $\rm{picosecond}/\sqrt{km}$. Let’s denote $\mathcal{P}_{a,sy}=\mathcal{P}_{a}+\mathcal{P}_{sy}$. The good fit of $\mathcal{P}_{a,sy}$ exhibits a small modulation on the exponential function of $e^{-2\mu\frac{c}{n_{r}}t}$ as shown (red line) in in Fig.3. The modulation is caused by the harmonic function of $\sqrt{t}$ in $\mathcal{P}_{a,sy}$. This implies that the probability $P$ of the Werner state exhibits non-Markovian dynamic. We substitute the experiment value of $\Delta\omega=2\pi\times 200\rm{GHz}$ in $\mathcal{P}_{a,sy}$. We obtain $D_{p1}=0.0017\,ps$ and $D_{p2}=0.047\,ps$. Then, the total PMD parameter $D_{p}=D_{p1}+D_{p2}=0.05\,ps$ which is in agreement with the PMD link value of $\leq 0.1\frac{ps}{\sqrt{km}}$ (Corning SMF-28 Optical Fiber data sheet). We also obtain the loss parameter $\mu=6.0\times 10^{-6}\,m^{-1}$ from the fitting of the experimental data. We can rewrite the loss parameter $\mu$ as 0.006/km or 0.025 dB/km. This value is below 0.05 dB/km which is the induced attenuation caused by the environmental effects such as temperature and humidity.
@@
We also fit the experimental data with the summation of the probability functions $p_{1}(t)$ of Eq.(8) and $p_{2}(t)$ of Eq.(9) in Methods. Let’s denote $p_{3}(t)=p_{1}(t)+p_{2}(t)$. The $p_{3}(t)$ exhibits stronger oscillation than $\mathcal{P}_{a,sy}$ within the error bar of the data as shown (blue line) in Fig.3. The equation of $p_{3}(t)$ indicates stronger non-Markovian dynamics than the $\mathcal{P}_{a,sy}$ because the harmonic function of $p_{3}(t)$ is a function of $t$. We obtain the parameters $\kappa_{1}=753\,{s}^{-1}$, $\kappa_{2}=3528\,{s}^{-1}$, and $\gamma_{\circ}=16292\,{s}^{-1}$ from the fitting of the experimental data. The total $\kappa=4281\,{s}^{-1}$. We have $4\times\kappa=17124\,{s}^{-1}$ greater than $\gamma_{\circ}$ satisfying the non-Markovian criterion i.e. $4\kappa\geq\gamma_{\circ}$. Our experimental data confirms the non-Markovian dynamic of the buffered qubit.
@@
We further prove the existence of the Markovian process if the idler photon is not buffered. We replace the buffer in the idler channel with a free-space attenuator and adjust the attenuation $\alpha$ such that it equals the propagation loss (0.2 dB/km) $\times$ the fiber length. We then perform the quantum state tomography on the signal and idler. We obtain the average probability $P$ by using the same approach as discussed above. We plot the probability $P$ (blue data point) and also fit the data with any arbitrary exponential function as shown (blue dotted line) as shown in Fig.3. This indicates that the idler in free-space configuration is experiencing the Markovian dynamic without the fiber buffer i.e. without the coupling to the fiber birefringence and the fiber reservoir.
@@
Since the probability functions $\mathcal{P}_{a,sy}$ and $p_{3}(t)$ fit well with the measured probability $P$ of the Werner state, we can use the $\mathcal{P}_{a,sy}$ and $p_{3}(t)$ for quantum mutual information to quantify the quantum and classical correlation of the buffered photon and its paired photon. For a bipartite system, the total correlation, which is a measure of quantum mutual information, is given by $\mathcal{I}(\rho^{f})=\mathcal{C}(\rho^{f})+\mathcal{Q}(\rho^{f})$, where the $\mathcal{C}(\rho^{f})$ is classical correlation and the $\mathcal{Q}(\rho^{f})$ is quantum discord. For the Werner state, the quantum discord is given by,
@@
$\mathcal{C}(\rho^{f})=\frac{(1-P)}{2}\rm{log}_{2}(1-P)+\frac{(1+P)}{2}\rm{log}_{2}(1+P)$ (7)
@@
Figure 4: The experimental data: quantum discord (blue triangle), classical (red triangle), and concurrence (orange circle). The quantum discord $\mathcal{Q}$ (blue line and blue dotted line), classical $\mathcal{C}$ (black line and black dotted line), and concurrence $\mathcal{C}_{n}$ (orange line and orange dotted line) of the Werner state as a function of buffer time by using the probability function $\mathcal{P}_{a,sy}$ and $p_{3}(t)$, respectively.
@@
We plot the quantum discord $\mathcal{Q}(\rho^{f})$, classical correlation $\mathcal{C}(\rho^{f})$, and concurrence $\mathcal{C}_{n}$ as a function of the buffer time as shown in Fig.4. The experimental data is calculated based on the Eq.(6) where the probability $P$ is obtained from the measured density matrix of the Werner state. We also calculate the quantum discord, classical, and concurrence by using the $P=\mathcal{P}_{a,sy}$ (solid line) and the $p_{3}(t)$ (dotted line) as shown in Fig.4. See Fig.S2 in the Supplementary for the extrapolation plot of the $\mathcal{I}(\rho^{f})$, $\mathcal{C}(\rho^{f})$, $\mathcal{Q}(\rho^{f})$ and concurrence $\mathcal{C}_{n}$ as a function of the buffer time up to 1.5 ms. When $P>\frac{1}{3}$ which is the well-known Werner’s criterion for inseparable, the concurrent is not zero, so the Werner state is still entangled or inseparableZurek_prl_2001 . When the $P=\frac{1}{3}$, the concurrent is 0. This occurs at the buffer time = 0.9 ms (190 km) which is numerically obtained from the fitting curve of the function $\mathcal{P}_{a,sy}=\frac{1}{3}$ (see Fig.S2 and Fig.S3(a) in the Supplementary). As for the function $p_{3}(t)=\frac{1}{3}$, this falls at the buffer time = 0.4 ms (80 km) (see Fig.S2 and Fig.S3(b) in the Supplementary). The strength of the non-Markovian effect can determine the buffer time (fiber length) where the buffered photon can still preserve quantum entanglement. Since the function $\mathcal{P}_{a,sy}$ is a better fit, our data indicates that the fiber-based quantum buffer can preserve quantum entanglement with a buffer time of up to 0.9 ms.
@@
Another interesting feature portrayed by the Werner state is the relationship between its concurrence, quantum discord, and classical correlation. For the function $\mathcal{P}_{a,sy}$ in the range $10.523<\mathcal{P}_{a,sy}<$ 1, the concurrence is larger than the quantum discord and classical correlation for the buffer time (length) from 0 s(0 m) to 0.7 ms(145 km). The $\mathcal{P}_{a,sy}=0.532$ occurs at the buffer time = 0.7 ms as shown in Fig.S2 and Fig.S3(a) in the Supplementary. Similarly, for the function $p_{3}(t)=0.532$, this falls at the buffer time = 0.3 ms (64 km) as shown in Fig.S2 and Fig.S3(b) in the Supplementary. The most important feature of the quantum mutual information is the quantum discord $\mathcal{Q}(\rho^{f})$ $>$ the concurrence and the classical correlation. This occurs at the buffer time larger than 0.7 ms(0.3 ms) for the function $\mathcal{P}_{a,sy}$ ($p_{3}(t)$), respectively.
@@ III Discussion
To our knowledge, this measurement is the first demonstration based on a simple qubit-fiber birefringence-reservoir model to study the non-Markovian effect on a two-photon polarization entangled state in which one photon from the pair is stored in a buffer made of different fiber lengths. We compare our model with the traditional atom-cavity-reservoir model. The function $\mathcal{P}_{a,sy}$ fits the experimental data better than the $p_{3}(t)$. This is because the harmonic function of $\mathcal{P}_{a,sy}$ is a function of $\sqrt{t}$ due to the PMD effect in our model. While the harmonic function of $p_{3}(t)$ is a function of $t$ (see Methods for atom-cavity-reservoir model). We first fit our experimental data with the $\mathcal{P}_{a,sy}$ which is the summation of Eq.(3) and Eq.(4). We find that the contribution from Eq.(4) is more dominant than the Eq.(3). This implies that the H and V polarization components of the buffered photon rotate in the same direction. This finding is a new insight into the polarization properties of the buffered qubit that undergoes the non-Markovian dynamic. We then fit our experimental data with the $p_{3}(t)$ which is the summation of Eq.(8) and Eq.(9). The obtained $\kappa$ is slightly above the threshold $\frac{\gamma_{\circ}}{4}$ which is the condition for an open quantum system to become non-Markovian. Since our quantum system is operating close to the threshold $\kappa\sim\frac{\gamma_{\circ}}{4}$, the configuration of our experiment presents the possibility of naturally occurring crossover between the Markovian and non-Markovian processes.
@@
We subtract the accidental coincidence in our data analysis so that the loss parameter $\mu$ is mainly contributed from the degradation of the polarization extinction ratio of the photonic qubit. One can regard the $\mu$ as the strength of the environment depolarization effect on the buffered photon. The obtained value of $\mu$ is below the induced attenuation value 0.05 dB/km caused by environmental factors such as temperature and humidity. Our finding indicates that the density matrix $\rho(t)$ of the entangled state at the output of the buffer is related to its early $\rho(t^{\prime})$ in the buffer, where $t>t^{\prime}$. This occurs when the environment depolarization effect of the fiber is more dominant than the PMD effect. We also demonstrate the Markovian scenario where the PMD effect did not exist by replacing the buffer with a free-space attenuator representing a memory-less environment (fiber birefringence and fiber reservoir).
We use the Werner state as the final state for the buffered idler photon and its paired signal photon. The Werner state is entangled (inseparable) when the $P>\frac{1}{3}$. The concurrence is zero at $P=\frac{1}{3}$ which is the Werner’s well-known separability criterion. The entanglement is related to the concurrence as $E_{n}(\rho_{AB})=H((1+\sqrt{1-\mathcal{C}_{n}^{2}})/2)$, where the $\mathcal{C}_{n}$ is the concurrence, and the function $H(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$ is the Shannon entropy. When the concurrence is zero, the entanglement is also zero at $P=\frac{1}{3}$. Based on this criterion, we can design the optimum fiber length of up to 190 km for the buffer in which the state is still inseparable or entangled. This buffer length corresponds to the buffer time of 0.9 ms.
@@
Several measures of detecting the non-Markovian dynamics have been proposedLuo_pra_2012 ; Plenio_PRL_2010 ; Piilo_16 ; Piilo_09 ; Piilo_10 . The most recent method uses quantum process capabilities of non-completely-positive (CP) processes Chan_24 for identifying the non-Markovian dynamics. Our approach here of using quantum state tomography to find the probability function $p(t)$ leads us to the measures of quantum mutual information such as classical correlation and quantum discord. We then use the quantum discord to characterize the quantum correlation of the entangled state in our buffer system. Quantum discord is a better concept than entanglement to characterize the quantum correlation of an open quantum system Zurek_prl_2001 . The quantum discord is nonzero at $P=\frac{1}{3}$ where the concurrence is zero. (see Fig.S3 in the Supplementary). The non-zero quantum discord can surpass Werner’s criterion on the optimum buffer time. The quantum discord implies that the quantum bi-partite correlation still exists for buffer time greater than 0.9 ms.
@@
The possibility for the buffered photon to experience the case $\Delta\phi_{h}\neq\Delta\phi_{v}$ is rare for the single-mode fiber which is used for communication networks. The reason is that the distributed fiber has a homogenous refractive index. The PMD effect of a light field in the single-mode fiber should not depend on the input polarization state of the light. However, the case $\Delta\phi_{h}\neq\Delta\phi_{v}$ may occur for the photon buffered in a loop of fiber more than one round trip. We are currently investigating the quantum buffer made of a fiber loop in a 2x2 Switch configurationkflee24 , where the photon is buffered by passing the same fiber loop multiple times.
Our analytical approach to derive the probability function can be easily extended to two qubits interacting separately with two independent non-Markovian environments such as the idler and the signal are buffered independently in two different buffers. It has been demonstrated that the non-Markovian density matrix evolution on one qubit state can mathematically proceed to the density matrix of a two-qubit entangled state Breuer_oxford_2002 ; Petruccione_PRA_2006 ; Bellomo_PRL_2007 ; Bellomo_PRA_2008 . The strength of the non-Markovian effect will be proportional to the square of Eq.(3) and Eq.(4), and hence will reduce the optimum buffer time. However, one can design the parameters $\kappa$ and $\gamma_{\circ}$ so that the crossover between Markovian and non-Markovian on the qubits can be controlled and utilized for modeling quantum networks Helsen_2023 ; Plenio_2014 and protecting quantum entanglementWenjie_2023 ; Franco_23 . We may improve the fiber-based quantum buffer by using the ultra-low loss single mode fiber such as SMF-28-ULL (Corning) and low loss Z-fiber from Sumitomo Electric, so that we can control the buffer to exhibit the Markovian effect, where its $\rho(t)$ is not related to $\rho(t^{\prime})$ in the buffer or the buffered qubit should not have memory effect on the previous/later qubit.
IV Methods

### v3

@@ II Results
We characterize the prepared state $|\psi_{\circ}\rangle=(|H_{s}H_{i}\rangle+|V_{s}V_{i}\rangle)/\sqrt{2}$ using the standard method of QST James_pra_2001 ; White_pra_2001 ; Munro_pra_2001 . There are a total of 16 settings for the HWPs and QWPs in the $\rm{PA_{s}}$ and $\rm{PA_{i}}$. For each setting, we measure the coincidence counts (CC) and accidentals (AC) with an integration time of 2 seconds or 100M sampling gates while the switches are turned on. Using maximum likelihood estimation, we reconstruct the $4\times 4$ density matrix of the final state $\rho^{f}$ in the HV basis. We construct the density matrix $\rho^{f}$ based on the true coincidence counts, i.e., the accidental coincidence counts are subtracted. In other words, noise photons generated via Raman scattering and loss-induced accidentals do not contribute to this study. We repeat the state tomography for different fiber lengths from 0.0 m up to 25 km. The 0.0 m fiber length corresponds to the short path in the buffer.
@@
We use quantum discord to characterize the non-Markovian dynamics of the buffered photon and its paired photon. It is of both fundamental and practical interest to differentiate classical and quantum correlations within the total correlation for the buffered photon experiencing a non-Markovian environment. Two scenarios can contribute to the classical information: (1) the environment projects or makes a measurement on the idler qubit state and creates the classical state, (2) the environment’s perturbation can dechore the information of the idler qubit state such as degradation of the polarization properties. Scenario (1) is the classical correlation that will appear at the coincidence and accidental coincidence. We can avoid this classical correlation by subtracting out the accidental coincidence coounts. Scenario (2) can create a mixture of product states. We study the quantum discord under scenario (2), which allow us to have a buffer time $>$ 0.9 ms. For this purpose, we can accurately treat the final state as the Werner state. The Werner state is the mixture of the maximally entangled state with the probability $P$ and the maximally mixed state, as given by Werner_pra_1989 ; Munro_pra_2001 ,
@@
where the $I$ is the identity matrix. We calculate the average probability $P=\sum^{n=6}_{i}p^{\prime}_{i}/n$, where $p^{\prime}_{1(2)}=4\rho^{\prime}_{11(44)}-1$, $p^{\prime}_{3(4)}=2\rho^{\prime}_{14(41)}$, and $p^{\prime}_{5(6)}=1-4\rho^{\prime}_{22(33)}$. The density matrix elements $\rho^{\prime}_{11}$, $\rho^{\prime}_{44}$, $\rho^{\prime}_{14}$, $\rho^{\prime}_{41}$, $\rho^{\prime}_{22}$, and $\rho^{\prime}_{33}$ are obtained from the final state. The measured average probability $P$ is the non-Markovian probability function p(t). ( See Methods for amplitude damping on one photon from the two-photon Werner state).
@@
We plot the average probability $P$ of the Werner state as a function of buffer time $t$, as shown (red solid circle) in Fig.3. We fit the experimental data with the summation of the probability functions $\mathcal{P}_{a}$ of Eq.(3) and $\mathcal{P}_{sy}$ of Eq.(4). With the substitution of $L=(c/n_{r})t$ and $\sqrt{\langle\tau^{2}\rangle}=D_{p}\,\sqrt{L}$, the $\Delta\phi_{1,2}$ in the harmonic functions of $\mathcal{P}_{a}$ and $\mathcal{P}_{sy}$ are then given by $\Delta\phi_{1,2}=\Delta\omega\,\sqrt{\langle\tau_{1,2}^{2}\rangle}=\Delta\omega\,D_{p1,p2}\,\sqrt{\frac{c\cdot t}{n_{r}}}$, respectively, where $D_{p1,p2}$ is the PMD parameter in units of $\rm{picoseconds}/\sqrt{km}$. Let’s denote $\mathcal{P}_{a,sy}=\mathcal{P}_{a}+\mathcal{P}_{sy}$. The good fit of $\mathcal{P}_{a,sy}$ exhibits a small modulation on the exponential function of $e^{-2\mu\frac{c}{n_{r}}t}$, as shown (red line) in Fig.3. The modulation is caused by the harmonic function of $\sqrt{t}$ in $\mathcal{P}_{a,sy}$. This implies that the probability $P$ of the Werner state exhibits non-Markovian dynamics. We substitute the experimental value of $\Delta\omega=2\pi\times 200\rm{GHz}$ in $\mathcal{P}_{a,sy}$. We obtain $D_{p1}=0.0017\,\frac{ps}{\sqrt{km}}$ and $D_{p2}=0.047\,\frac{ps}{\sqrt{km}}$. The $D_{p2}\,>\,D_{p1}$ indicates that the harmonic function of Eq.(4) is more dominant than Eq.(3). Then, the total PMD parameter $D_{p}=D_{p1}+D_{p2}=0.05\,\frac{ps}{\sqrt{km}}$, which is in agreement with the PMD link value of $\leq 0.1\frac{ps}{\sqrt{km}}$ (Corning SMF-28 Optical Fiber data sheet). We also obtain the loss parameter $\mu=6.0\times 10^{-6}\,m^{-1}$ from the fitting of the experimental data. We can rewrite the loss parameter $\mu$ as 0.006/km or 0.025 dB/km. This value is below 0.05 dB/km, which is the induced attenuation caused by environmental effects such as temperature and humidity (Corning SMF-28 Optical Fiber data sheet).
@@
We also fit the experimental data with the summation of the probability functions $p_{1}(t)$ of Eq.(8) and $p_{2}(t)$ of Eq.(9) in the Methods section. Let’s denote $p_{3}(t)=p_{1}(t)+p_{2}(t)$. The $p_{3}(t)$ exhibits stronger oscillation than $\mathcal{P}_{a,sy}$ within the error bar of the data, as shown (blue line) in Fig.3. The equation of $p_{3}(t)$ indicates stronger non-Markovian dynamics than the $\mathcal{P}_{a,sy}$ because the harmonic function of $p_{3}(t)$ is a function of $t$. We obtain the parameters $\kappa_{1}=753\,{s}^{-1}$, $\kappa_{2}=3528\,{s}^{-1}$, and $\gamma_{\circ}=16292\,{s}^{-1}$ from the fitting of the experimental data. The $\kappa_{2}\,>\,\kappa_{1}$ indicates that the harmonic function of Eq.(9) is more dominant than Eq.(8). The total $\kappa=4281\,{s}^{-1}$. We have $4\times\kappa=17124\,{s}^{-1}$, which is greater than $\gamma_{\circ}$, satisfying the non-Markovian criterion, i.e., $4\kappa\geq\gamma_{\circ}$. Our experimental data confirms the non-Markovian dynamics of the buffered qubit and its paired photon.
@@
We further prove the existence of the Markovian process if the idler photon is not buffered. We replace the buffer in the idler channel with a free-space attenuator and adjust the attenuation $\alpha$ such that it equals the propagation loss (0.2 dB/km) $\times$ the fiber length. We then perform the quantum state tomography on the signal and idler. We obtain the average probability $P$ using the same approach as discussed above. We plot the probability $P$ (blue data point) and also fit the data with an arbitrary exponential function as shown (blue dotted line) as shown in Fig.3. This indicates that the idler in the free-space configuration is experiencing a Markovian environment without the fiber buffer, i.e., without the coupling to the fiber birefringence and the fiber reservoir.
@@
Since the probability functions $\mathcal{P}_{a,sy}$ and $p_{3}(t)$ fit well with the measured probability $P$ of the Werner state, we can use $\mathcal{P}_{a,sy}$ and $p_{3}(t)$ for quantum mutual information to quantify the quantum and classical correlation of the buffered photon and its paired photon. For a bipartite system, the total correlation, which is a measure of quantum mutual information, is given by $\mathcal{I}(\rho^{f})=\mathcal{C}(\rho^{f})+\mathcal{Q}(\rho^{f})$, where $\mathcal{C}(\rho^{f})$ is the classical correlation and $\mathcal{Q}(\rho^{f})$ is the quantum discord. For the Werner state, the quantum discord is given by,
@@
$\mathcal{C}(\rho^{f})=\frac{(1-P)}{2}\rm{log}_{2}(1-P)+\frac{(1+P)}{2}\rm{log}_{2}(1+P).$ (7)
@@
Figure 4: The experimental data: quantum discord (blue triangle), classical (red triangle), and concurrence (orange solid circle). The quantum discord $\mathcal{Q}$ (blue line and blue dotted line), classical $\mathcal{C}$ (black line and black dotted line), and concurrence $\mathcal{C}_{n}$ (orange line and orange dotted line) of the Werner state as a function of buffer time by using the probability function $\mathcal{P}_{a,sy}$ and $p_{3}(t)$, respectively.
@@
We plot the quantum discord $\mathcal{Q}(\rho^{f})$, classical correlation $\mathcal{C}(\rho^{f})$, and concurrence $\mathcal{C}_{n}$ as a function of the buffer time, as shown in Fig.4. The experimental data is calculated based on Eq.(6) where the probability $P$ is obtained from the measured density matrix of the Werner state. We also calculate the quantum discord, classical, and concurrence using the $P=\mathcal{P}_{a,sy}$ (solid line) and the $p_{3}(t)$ (dotted line) as shown in Fig.4. See Fig.S2 in the Supplementary for the extrapolation plot of the $\mathcal{I}(\rho^{f})$, $\mathcal{C}(\rho^{f})$, $\mathcal{Q}(\rho^{f})$ and concurrence $\mathcal{C}_{n}$ as a function of the buffer time up to 1.5 ms. When $P>\frac{1}{3}$, which is the well-known Werner’s criterion for inseparability, the concurrence is not zero, indicating that the Werner state is still entangled or inseparableZurek_prl_2001 . When the $P=\frac{1}{3}$, the concurrence is 0. This occurs at the buffer time of 0.9 ms (190 km), which is numerically obtained from the fitting curve of the function $\mathcal{P}_{a,sy}=\frac{1}{3}$ (see Fig.S2 and Fig.S3(a) in the Supplementary). As for the function $p_{3}(t)=\frac{1}{3}$, this occurs at the buffer time of 0.4 ms (80 km) (see Fig.S2 and Fig.S3(b) in the Supplementary). The strength of the non-Markovian effect can determine the buffer time (fiber length) where the buffered photon and its paired photon can still preserve quantum entanglement. Since the function $\mathcal{P}_{a,sy}$ is a better fit, our data indicates that the fiber-based quantum buffer can preserve quantum entanglement with a buffer time of up to 0.9 ms.
@@
Another interesting feature portrayed by the Werner state is the relationship between its concurrence, quantum discord, and classical correlation. For the function $\mathcal{P}_{a,sy}$ in the range $10.523<\mathcal{P}_{a,sy}<$ 1, the concurrence is larger than the quantum discord and classical correlation for the buffer time (length) from 0.0 s(0.0 m) to 0.7 ms(145 km). The $\mathcal{P}_{a,sy}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.523}$ occurs at the buffer time of 0.7 ms, as shown in Fig.S2 and Fig.S3(a) in the Supplementary. Similarly, for the function $p_{3}(t)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.523}$, this occurs at the buffer time of 0.3 ms (64 km), as shown in Fig.S2 and Fig.S3(b) in the Supplementary. The most important feature of the quantum mutual information is that the quantum discord $\mathcal{Q}(\rho^{f})$ is greater than the concurrence and the classical correlation. This occurs at the buffer time greater than 0.7 ms(0.3 ms) for the function $\mathcal{P}_{a,sy}$ ($p_{3}(t)$) $\leq$ 0.523, respectively.
@@ III Discussion
To our knowledge, this measurement is the first demonstration based on a simple qubit-fiber birefringence-reservoir model to study the non-Markovian effect on a two-photon polarization entangled state in which one photon from the pair is stored in a buffer made of different fiber lengths. We compare our model with the traditional atom-cavity-reservoir model. The function $\mathcal{P}_{a,sy}$ fits the experimental data better than $p_{3}(t)$. This is because the harmonic function of $\mathcal{P}_{a,sy}$ is a function of $\sqrt{t}$ due to the PMD effect in our model, while the harmonic function of $p_{3}(t)$ is a function of $t$ (see Methods for atom-cavity-reservoir model). We find that the contribution from Eq.(4) is more dominant than Eq.(3). This implies that the H and V polarization components of the buffered photon rotate in the same direction. This finding provides new insight into the polarization properties of the buffered qubit in a non-Markovian environment. Since our quantum system is operating close to the threshold $\kappa\sim\frac{\gamma_{\circ}}{4}$, the configuration of our experiment presents the possibility of naturally occurring crossover between the Markovian and non-Markovian processes.
@@
Our finding indicates that the density matrix $\rho(t)$ of the entangled state at the output of the buffer is related to its early state $\rho(t^{\prime})$ in the buffer, where $t>t^{\prime}$. This occurs when the environment depolarization effect of the fiber is more dominant than the PMD effect. We also demonstrate the Markovian scenario where the PMD effect did not exist by replacing the buffer with a free-space attenuator.
@@
Several measures of detecting the non-Markovian dynamics have been proposed Luo_pra_2012 ; Plenio_PRL_2010 ; Piilo_16 ; Piilo_09 ; Piilo_10 . The most recent method uses quantum process capabilities of non-completely-positive (CP) processes Chan_24 for identifying the non-Markovian dynamics. Our approach here uses quantum state tomography to find the probability function $p(t)$, which leads us to the measures of quantum mutual information such as classical correlation and quantum discord.
@@
The possibility for the buffered photon to experience the case $\Delta\phi_{h}\neq\Delta\phi_{v}$ is rare for the single-mode fiber used in communication networks. The reason is that the distributed fiber has a homogenous refractive index. The PMD effect of a light field in the single-mode fiber should not depend on the input polarization state of the light. However, the case $\Delta\phi_{h}\neq\Delta\phi_{v}$ may occur for the photon buffered in a loop of fiber more than one round trip. We are currently investigating the quantum buffer made of a fiber loop in a 2x2 switch configuration kflee24 , where the photon is buffered by passing through the same fiber loop multiple times.
IV Conclusion
We have introduced a qubit-fiber-reservoir model to explore the non-Markovian dynamics of the two-photon polarization-entangled state. The non-Markovian probability function obtained from our approach fits well with the experimental data. One can design the model’s parameters so that the crossover between Markovian and non-Markovian dynamics on the qubits can be controlled and utilized for modeling quantum networks Helsen_2023 ; Plenio_2014 and protecting quantum entanglement Wenjie_2023 ; Franco_23 . Our analytical approach to deriving the probability function can be easily extended to two qubits interacting separately with two independent non-Markovian environments, with the idler and the signal buffered independently in two different buffers Breuer_oxford_2002 ; Petruccione_PRA_2006 ; Bellomo_PRL_2007 ; Bellomo_PRA_2008 . The non-Markovian effect will be strong, hence reducing the optimum buffer time. Our approach provides new insight into the polarization properties of the buffered photon. This will be very beneficial for developing variable quantum buffers for controlling packet delays and allowing flexible contention management in quantum networks using classical-quantum data frames Shabani_22 ; BenYoo_2024 and a sequence of ordered signals Giovannetti_JPA_05 . Our model can be easily adapted into the formalism of tensor network states Lubasch_18 to model quantum processing.
V Methods